# Global existence and boundedness in an attraction-repulsion chemotaxis system with nonlocal logistic source and sublinear productions

### v1

@@ Abstract
in an open, bounded domain $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ with smooth boundary $\partial\Omega$. Assume the parameters $d_{1}$, $d_{2}$, $d_{3}$, $\chi$, $\xi$, $\alpha$, $\beta$ and $\mu$ are positive constants, initial data $({u}_{0},{v}_{0},{w}_{0})$ are nonnegative and the function $f(u)\leq Ku^{l}\in C^{1}([0,\infty))$ for some $K,l>0$. Under appropriate conditions on the parameter $k$, $l$ and $m$ we show that the above problem admits a unique globally bounded classical solution. Further, to illustrate the analytical results, some of numerical examples are provided.
@@
Attraction-repulsion model, Chemotaxis system, Classical solution, Global existence, Non-local source, Numerical simulation
@@ 1 Introduction and motivation
$\left\{\begin{array}[]{rrll}&&{u}_{t}=d_{1}\Delta{u}-\chi\nabla\cdot({u}^{k}{\nabla{v}})+\xi\nabla\cdot({u}^{k}{\nabla{w}})+\displaystyle{\mu{u}^{m}\left(1-\int_{\Omega}{u}(x,t)\mathrm{d}x\right)},&x\in\Omega,\,t>0,\\ &&{v}_{t}=d_{2}\Delta{v}-\alpha{v}+f(u),&x\in\Omega,\,t>0,\\ &&{w}_{t}=d_{3}\Delta{w}-\beta{w}+f(u),&x\in\Omega,\,t>0,\\ &&\frac{\partial{u}}{\partial\nu}=\frac{\partial{v}}{\partial\nu}=\frac{\partial{w}}{\partial\nu}=0,&x\in\partial\Omega,\,t>0,\\ &&{u}(x,0)={u}_{0},\quad{v}(x,0)={v}_{0},\quad{w}(x,0)={w}_{0},&x\in\Omega,\end{array}\right.$ (1.6)
in an open, bounded domain $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ with smooth boundary $\partial\Omega$ and here $\nu$ denotes the outward unit normal on $\partial\Omega$. We assume the parameters $d_{1}$, $d_{2}$, $d_{3}$, $\chi$, $\xi$, $\alpha$, $\beta$ and $\mu$ are positive constants and the initial data $({u}_{0},{v}_{0},{w}_{0})\in{{C^{0}}(\overline{\Omega})}\times{{W^{1,q}}(\overline{\Omega})}\times{{W^{1,q}}(\overline{\Omega})}$ are nonnegative. The unknown functions ${u}$, ${v}$ and ${w}$ represent the cell density, the concentration of attractive and repulsive chemical substances, respectively. The system (1.6) describes the interaction of one cell with two different chemicals secreted by the cells. The attraction effect is given by the term $-\chi\nabla\cdot(u^{k}\nabla v)$, where $\chi$ is the attraction coefficient. On the other hand, the term $\xi\nabla\cdot(u^{k}\nabla w)$ represent the repulsion effect, where $\xi$ is the repulsion coefficient. The proliferation rate of cells are assumed to have a logistic growth in the form of nonlocal source and is given by $\mu u^{m}(1-\int_{\Omega}u)$, where $\mu$ is the growth rate constant. Here, the decay rates of the chemicals are given by $\alpha v$ and $\beta w$ respectively. Here the growth rate of the cells is named by the function $f(u)$. Further, $d_{1}$, $d_{2}$ and $d_{3}$ are positive diffusion coefficients of cell density and chemical signals respectively.
@@
$\left\{\begin{array}[]{rrll}{u}_{t}&=&\Delta{u}-\nabla\cdot({u}\nabla{v}),\\ {v}_{t}&=&\Delta{v}-{v}+{u},\end{array}\right.$ (1.13)
where ${u}$ and ${v}$ denote the cell density and the concentration of chemical signals, respectively. Analytic and numerical simulations of various Keller-Segel models and Keller-Segel models with logistic source have been studied in the past few decades. For the recent evolution in this field one can refer the articles by Lankeit and Winkler [24], Bellomo et al. [1] , Horstmann [17] and the references therein.
To the better understanding of the system (1.6), Luca [33] propsed the attraction-repulsion system
@@
For the parabolic-elliptic case, where $D({u})\geq C_{D}({u}+\zeta)^{m-1}$, $m\geq 1$ and $f({u})\leq a-b{u}^{\eta}$, $\eta>1$ Wang [39] showed the existence of global bounded classical solutions to the system (1.21) for the case $\zeta>0$, and also, when either $\xi\gamma-\chi\alpha>0$ or logistic damping is large or diffusion is sufficiently large enough. Further, the system has weak solutions for the case $\zeta=0$ under the same assumptions. In addition, the asymptotic behavior has been studied to the system (1.21) with particular logistic source. Under some suitable conditions on the logistic source and the parameters, Zhang et. al. [48] obtained the global bounded classical solution as well as the weak solution to the system (1.21) with $D({u})\geq-C{u}(1+{u})$ and the large time behavior for $f({u})=\mu{u}(1-{u})$ respectively. Zhao et. al. [49] investigated the global bounded solution when $f({u})=r{u}-\mu{u}^{2}$ under the condition that $\chi\alpha-\xi\gamma>0$ and $\mu=\frac{1}{3}(\chi\alpha-\xi\gamma)$ with $D(u)\geq C_{D}u^{m-1}$, $m\geq 1$. Further, the exponentially convergence result is provided for the case when $D({u})=1$ and $n\geq 3$. On the other hand, they studied the boundedness and blow-up for the case when $D({u})=0$ and $n\geq 2$. When $D({u})\geq C_{D}({u}+1)^{m-1}$, $m\geq 1$ and $f({u})\leq a-b{u}^{\eta}$, $\eta>1$, Yan et. al. [45] proved the global bounded solution to the system (1.21) and this solution may blow-up whenever $\|{u}_{0}\|_{{\mathcal{L}^{1}}(\Omega)}>\frac{8\pi}{\chi\alpha-\xi\gamma}$ and $\chi\alpha-\xi\gamma>0$ under the specific functions in two dimensional settings. For the recent article see, [10].
@@
Li et. al. [27] proved the system (1.21) has global bounded solutions for both the cases $\tau=0,1$ with $f({u})\leq a-b{u}^{\rho}$. For the case $\tau=0$ and $\xi\gamma-\chi\alpha>0$, there exits a bounded classical solution to the system (1.21) whenever $\rho\geq 1$. On the other hand, if $\xi\gamma-\chi\alpha<0$, still the result is valid whenever $\rho$ is strong enough. In one and two space dimensions, the author established the same results when $\tau>0$. For the fully parabolic case, Zheng et. al. [50] proved the existence of global bounded classical solution to the system (1.21) with $f({u})\leq a{u}-\mu{u}^{2}$ under some suitable conditions and this result improved the result of Li et. al. [27] provided $n\geq 3$ and $\beta=\delta$. In three dimensional settings, for $\beta,\delta\geq\frac{1}{2}$ and $\mu\geq\max\{(\frac{41}{2}\chi\alpha+9\xi\gamma)^{\eta},(9\chi\alpha+\frac{41}{2}\xi\gamma)^{\eta}\}$, the global bounded classical solutions to (1.21) with $f({u})={u}-\mu{u}^{\eta+1}$, $\eta>1$ has been studied by Li et. al. [25] and this solution converges to the steady state as $t\to\infty$. If, $\chi\alpha=\xi\gamma$, Shi et. al. [38] established the global bounded classical solutions to the system (1.21) with $f({u})\leq a-b{u}^{\theta}$ under $\theta$ and $n$ satisfying some suitable conditions. If $\mu>\frac{\chi^{2}\alpha^{2}(\beta-\delta)^{2}}{8\delta\beta^{2}}$, the exponential stability is discussed with $f({u})=\mu{u}(1-{u})$ for $n\leq 9$ by constructing appropriate Lyapunov functional. Suppose $D({u})\geq C{u}^{m-1}$ and also $m>\frac{6}{5}$, the global bounded classical solution is obtained by Zeng [47]. On the other hand, same results holds for $D({u})\leq C({u}^{m-1}+1)$ when $m\in(1,\frac{6}{5}]$. For more recent articles, one can refer [6, 14, 36].
@@
The existence of global solution is proved under some conditions on the attraction with max$\{l,s,\frac{2}{n}\}>k$. They also showed that the size of the coefficients will affect the boundedness of the solution and these results improved the results of [48]. Liu et. al. [32] established the finite time blowup results to the following system with nonlinear production
@@
If $\alpha\geq 1$, $m>1$, $\gamma\geq 1$, $\alpha+a>m+\gamma$, $a_{1}>0$, $a_{1}-a_{2}|\Omega|>0$ and under some suitable conditions on the initial data, the authors proved the global existence of solutions to the considered system. Moreover, the solution converges to the stead state ${u^{*}=\frac{a_{0}^{\frac{1}{\alpha}}}{(a_{1}-a_{2}|\Omega|)^{\frac{1}{\alpha}}},v^{*}=(u^{*})^{\gamma}}$.
@@
for $\tau=1$ and $\alpha,\beta\geq 1$. The authors established the uniform boundedness of classical solutions under two key regimes: the subquadratic case, $21\leq\alpha<$ 2 and $\beta>\frac{n+4}{2}-\alpha$; the superquadratic case, $\beta>\frac{n}{2},2\leq\alpha<1+\frac{2\beta}{n}$. This work extends the analysis of the fully parabolic case previously studied in [4].
Funtes et al. [12] studied the following system
@@
Motivated by the above research works, main purpose of this paper is to establish the existence of the global classical solution to the system (1.6) with suitable conditions on $k$, $l$ and $m$, which is uniformly bounded in $\Omega\times(0,\infty)$ for $n\geq 2$. More to the point, this study is to examine the simultaneous effects of both attraction and repulsion mechanism in the presence of nonlocal source and sublinear production terms. Moreover, some specific comparisons are made for different values of parameters in numerical simulations.
Our article is organized as follows: Section 2, contains some basic inequalities, lemmas and the local existence of classical solution. Section 3 deals with boundedness and global existence of classical solution to the system (1.6). Section 4 is devoted to the presentation and discussion of our numerical examples. Finally, section 5 gives the conclusion of our work.
@@
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ is an open, bounded domain with smooth boundary and $q>n$. Moreover, the function $f$ fulfills (1.10) with $l+k<1+\frac{2}{n}$, where $k\geq 1$ and $1<m<1+\frac{2}{n}$. Assume that, there exists $p>\frac{n}{2}$, then for any initial data $({u}_{0},{v}_{0},{w}_{0})$ satisfy (1.9), then the system (1.6) possesses a unique classical solution $({u},{v},{w})$ which is uniformly bounded in the sense that
@@
Remark 1.1.
For the case $n=1$, the solution (u, v, w) exists globally by Lemma 2.4 with $p=1$, without any additional restrictions on the parameters.
@@ 2 Preliminaries and local existence
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 2,$ is an open, bounded domain with smooth boundary and assume that $q>n$ and $M_{0}>0$. Moreover, the function $f$ fulfills (1.10) and $k,m\geq 1$, then for each non-negative initial data satisfies (1.9), there exists $T_{\mathrm{max}}\in(0,\infty]$ such that a uniquely determined triple $({u},{v},{w})$ of nonnegative functions
@@
The proof can be derived by the standard arguments involving the parabolic regularity theory. For the proof we refer to [18] and [11]. In addition, the non-negativity of the solution in $\Omega\times(0,T_{\mathrm{max}})$ is comes from the maximum principle along with (1.9). ∎
@@
(See [3]) For $\int_{\Omega}{u}(\cdot,0)=M_{0}>0$, the mass $\int_{\Omega}{u}(x,t)\mathrm{d}x=M(t)$ satisfies
@@
For each $y_{0}\in{\mathcal{W}^{2,r}}(\Omega)$ such that $\frac{\partial y(\cdot,s_{0})}{\partial\nu}=0$ on $\partial\Omega$ and any $g\in{\mathcal{L}^{r}}\big{(}(0,T);{\mathcal{L}^{r}}(\Omega)\big{)}$, there exists a unique solution
@@
$\int_{s_{0}}^{T}\int_{\Omega}y^{r}+\int_{s_{0}}^{T}\int_{\Omega}|y_{t}|^{r}+\int_{s_{0}}^{T}\int_{\Omega}|\Delta y|^{r}\leq C_{r}\int_{s_{0}}^{T}\int_{\Omega}|g|^{r}+C_{r}\int_{\Omega}y_{0}^{r}+C_{r}\int_{\Omega}|\Delta y_{0}|^{r}.$
@@
$\int_{s_{0}}^{T}\int_{\Omega}e^{sr}|\Delta y|^{r}\leq C_{r}\int_{s_{0}}^{T}\int_{\Omega}e^{sr}|g|^{r}+C_{r}\int_{\Omega}y(\cdot,s_{0})^{r}+C_{r}\int_{\Omega}|\Delta y(\cdot,s_{0})|^{r}.$
@@
Let $n\geq 1$, $M_{0}>0$ and the initial data ${u}_{0}$, ${v}_{0}$ and ${w}_{0}$ satisfy (1.9) for some $q>n$. Moreover, the function g fulfill (1.10) with $l+k<1+\frac{2}{n}$, where $k\geq 1$ and $1<m<1+\frac{2}{n}$. Suppose that there exists $p\geq 1$ such that $p>\frac{n}{2}$ and
@@
$\sup\limits_{t\in(s_{0},T_{\mathrm{max}})}\Big{(}\big{\|}{u}(\cdot,t)\big{\|}_{{\mathcal{L}^{\infty}}(\Omega)}+\big{\|}{v}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}+\big{\|}{w}(\cdot,t)\big{\|}_{{\mathcal{W}^{1,q}}(\Omega)}\Big{)}<\infty.$ (2.2)
@@
$\frac{np}{(n-p)_{+}}=\left\{\begin{array}[]{rrll}&&\infty,&\quad\mbox{if}\quad p\geq n,\\ \\ &&\frac{np}{n-p}>n,&\quad\mbox{if}\quad\frac{n}{2}<p<n,\end{array}\right.$ (2.6)
[…]
${v}(\cdot,t)=e^{-\alpha t}e^{d_{2}t\Delta}{v}_{0}+\int^{t}_{0}e^{-\alpha(t-s)}e^{d_{2}(t-s)\Delta}f\big{(}u(\cdot,t)\big{)}\mathrm{d}s.$
@@
$+K\int^{t}_{0}e^{-\alpha(t-s)}\Big{\|}\nabla e^{d_{2}(t-s)\Delta}{u}^{l}(\cdot,s)\Big{\|}_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}\mathrm{d}s.$

### v2

@@ Abstract
in a bounded domain $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ with smooth boundary $\partial\Omega$. Assume the parameters $d_{1}$, $d_{2}$, $d_{3}$, $\chi$, $\xi$, $\alpha$, $\beta$ and $\mu$ are positive constants and $k,m\geq 1$. The initial data $({u}_{0},{v}_{0},{w}_{0})$ are nonnegative and the function $f(u)\leq Ku^{l}\in C^{1}([0,\infty))$ for some $K,l>0$. Under appropriate conditions on the parameter $k$, $l$ and $m$ we show that the above problem admits a unique globally bounded classical solution.
@@
Attraction-repulsion model, Chemotaxis system, Classical solution, Global existence, Non-local source
@@ 1 Introduction and motivation
$\left\{\begin{array}[]{rrll}&&{u}_{t}=d_{1}\Delta{u}-\chi\nabla\cdot({u}^{k}{\nabla{v}})+\xi\nabla\cdot({u}^{k}{\nabla{w}})+\displaystyle{\mu{u}^{m}\left(1-\int_{\Omega}{u}(x,t)\mathrm{d}x\right)},&x\in\Omega,\,t>0,\\ &&{v}_{t}=d_{2}\Delta{v}-\alpha{v}+f(u),&x\in\Omega,\,t>0,\\ &&{w}_{t}=d_{3}\Delta{w}-\beta{w}+f(u),&x\in\Omega,\,t>0,\\ &&\frac{\partial{u}}{\partial\nu}=\frac{\partial{v}}{\partial\nu}=\frac{\partial{w}}{\partial\nu}=0,&x\in\partial\Omega,\,t>0,\\ &&{u}(x,0)={u}_{0},\qquad{v}(x,0)={v}_{0},\qquad{w}(x,0)={w}_{0},&x\in\Omega,\end{array}\right$. (1.6)
in a bounded domain $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ with smooth boundary $\partial\Omega$ and here $\nu$ denotes the outward unit normal on $\partial\Omega$. We assume the parameters $d_{1}$, $d_{2}$, $d_{3}$, $\chi$, $\xi$, $\alpha$, $\beta$, $\mu$ are positive constants, $k,m\geq 1$ and the initial data are sufficiently regular. The unknown functions u{u}, v{v} and w{w} represent the cell density, the concentration of attractive and repulsive chemical substances, respectively. The system (1.6) describes the interaction of one cell with two different chemicals secreted by the cells. The attraction effect is given by the term $\chi\nabla\cdot(u^{k}\nabla v)$, where $\chi$ is the attraction coefficient. On the other hand, the term $\xi\nabla\cdot(u^{k}\nabla w)$ represent the repulsion effect, where $\xi$ is the repulsion coefficient. The proliferation rate of cells are assumed to have a logistic growth in the form of nonlocal source and is given by $\mu u^{m}(1-\int_{\Omega}u)$, where $\mu$ is the growth rate constant. Here, the decay rates of the chemicals are given by $\alpha v$ and $\beta w$ respectively. Here the growth rate of the cells is named by the function $f(u)$. Further, $d_{1}$, $d_{2}$ and $d_{3}$ are positive diffusion coefficients of cell density and chemical signals respectively.
@@
$\left\{\begin{array}[]{rrll}{u}_{t}&=&\Delta{u}-\nabla\cdot({u}\nabla{v}),\\ {v}_{t}&=&\Delta{v}-{v}+{u},\end{array}\right$.
where u{u} and v{v} denote the cell density and the concentration of chemical signals, respectively. Analytic and numerical simulations of various Keller-Segel models and Keller-Segel models with logistic source have been studied in the past few decades. For the recent evolution in this field one can refer the articles by Lankeit and Winkler [25], Bellomo et al. [1] , Horstmann [18] and the references therein.
To the better understanding of the system (1.6), Luca [34] proposed the attraction-repulsion system
@@
For the parabolic-elliptic case, where $D({u})\geq C_{D}({u}+\zeta)^{m-1}$, $m\geq 1$ and $f({u})\leq a-b{u}^{\eta}$, $\eta>1$ Wang [40] showed the existence of global bounded classical solutions to the system (1.18) for the case $\zeta>0$, and also, when either $\xi\gamma-\chi\alpha>0$ or logistic damping is large or diffusion is sufficiently large enough. Further, the system has weak solutions for the case $\zeta=0$ under the same assumptions. In addition, the asymptotic behavior has been studied to the system (1.18) with particular logistic source. Under some suitable conditions on the logistic source and the parameters, Zhang et. al. [49] obtained the global bounded classical solution as well as the weak solution to the system (1.18) with $D({u})\geq-C{u}(1+{u})$ and the large time behavior for $f({u})=\mu{u}(1-{u})$ respectively. Zhao et. al. [50] investigated the global bounded solution when $f({u})=r{u}-\mu{u}^{2}$ under the condition that $\chi\alpha-\xi\gamma>0$ and $\mu=\frac{1}{3}(\chi\alpha-\xi\gamma)$ with $D(u)\geq C_{D}u^{m-1}$, $m\geq 1$. Further, the exponentially convergence result is provided for the case when D(u)=1D({u})=1 and $n\geq 3$. On the other hand, they studied the boundedness and blow-up for the case when D(u)=0D({u})=0 and $n\geq 2$. When $D({u})\geq C_{D}({u}+1)^{m-1}$, $m\geq 1$ and $f({u})\leq a-b{u}^{\eta}$, $\eta>1$, Yan et. al. [46] proved the global bounded solution to the system (1.18) and this solution may blow-up whenever $\|{u}_{0}\|_{{\mathcal{L}^{1}}(\Omega)}>\frac{8\pi}{\chi\alpha-\xi\gamma}$ and $\chi\alpha-\xi\gamma>0$ under the specific functions in two dimensional settings. For the recent article see, [10].
@@
Li et. al. [28] proved the system (1.18) has global bounded solutions for both the cases $\tau=0,1$ with $f({u})\leq a-b{u}^{\rho}$. For the case $\tau=0$ and $\xi\gamma-\chi\alpha>0$, there exists a bounded classical solution to the system (1.18) whenever $\rho\geq 1$. On the other hand, if $\xi\gamma-\chi\alpha<0$, still the result is valid whenever $\rho$ is strong enough. In one and two space dimensions, the author established the same results when $\tau>0$. For the fully parabolic case, Zheng et. al. [51] proved the existence of global bounded classical solution to the system (1.18) with $f({u})\leq a{u}-\mu{u}^{2}$ under some suitable conditions and this result improved the result of Li et. al. [28] provided $n\geq 3$ and $\beta=\delta$. In three dimensional settings, for $\beta,\delta\geq\frac{1}{2}$ and $\mu\geq\max\{(\frac{41}{2}\chi\alpha+9\xi\gamma)^{\eta},(9\chi\alpha+\frac{41}{2}\xi\gamma)^{\eta}\}$, the global bounded classical solutions to (1.18) with $f({u})={u}-\mu{u}^{\eta+1}$, $\eta>1$ has been studied by Li et. al. [26] and this solution converges to the steady state as $t\to\infty$. If, $\chi\alpha=\xi\gamma$, Shi et. al. [39] established the global bounded classical solutions to the system (1.18) with $f({u})\leq a-b{u}^{\theta}$ under $\theta$ and $n$ satisfying some suitable conditions. If $\mu>\frac{\chi^{2}\alpha^{2}(\beta-\delta)^{2}}{8\delta\beta^{2}}$, the exponential stability is discussed with $f({u})=\mu{u}(1-{u})$ for $n\leq 9$ by constructing appropriate Lyapunov functional. Suppose $D({u})\geq C{u}^{m-1}$ and also $m>\frac{6}{5}$, the global bounded classical solution is obtained by Zeng [48]. On the other hand, same result holds for $D({u})\leq C({u}^{m-1}+1)$ when $m\in(1,\frac{6}{5}]$. For more recent articles, one can refer [6, 15, 37].
@@
[…]
@@
If $\alpha\geq 1$, $m>1$, $\gamma\geq 1$, $\alpha+1>m+\gamma$, $a_{1}>0$, $a_{1}-a_{2}|\Omega|>0$ and under some suitable conditions on the initial data, the authors proved the global existence of solutions to the considered system. Moreover, the solution converges to the stead state ${u^{*}=\frac{a_{0}^{\frac{1}{\alpha}}}{(a_{1}-a_{2}|\Omega|)^{\frac{1}{\alpha}}},v^{*}=(u^{*})^{\gamma}}$.
@@
for $\alpha,\beta\geq 1$. The authors established the uniform boundedness of classical solutions under two key regimes: the subquadratic case, $1\leq\alpha<2$ and $\beta>\frac{n+4}{2}-\alpha$; the superquadratic case, $\beta>\frac{n}{2},2\leq\alpha<1+\frac{2\beta}{n}$. This work extends the analysis of the fully parabolic case previously studied in [4].
Fuentes et al. [13] studied the following system
@@
Motivated by the above research works, main purpose of this paper is to establish the existence of the global classical solution to the system (1.6) with suitable conditions on $k$, $l$ and $m$, which is uniformly bounded in $\Omega\times(0,\infty)$ for $n\geq 2$. More to the point, this study is to examine the simultaneous effects of both attraction and repulsion mechanism in the presence of nonlocal source and sublinear production terms.
Our article is organized as follows: Section 2 contains some basic inequalities, lemmas and the local existence of classical solution. Section 3 deals with boundedness and global existence of classical solution to the system (1.6). Finally, the work ends with conclusion in Section 4.
@@
Suppose that $\Omega\subset\mathbb{R}^{n}\>(n\geq 2)$, is a bounded domain with smooth boundary and $q>n$. Moreover, the function $f$ fulfills (1.10) with $l+k<1+\frac{2}{n}$, where $k\geq 1$ and $1<m<1+\frac{2}{n}$. Then for any initial data $({u}_{0},{v}_{0},{w}_{0})$ satisfy (1.9), the system (1.6) possesses a unique classical solution (u,v,w)({u},{v},{w}) which is global and uniformly bounded in the sense that
@@ 2 Preliminaries and local existence
Definition 2.1 (Cauchy’s inequality with $\epsilon$ [11]).
$ab\leq\epsilon a^{2}+\frac{1}{4\epsilon}b^{2},\qquad\qquad a,b>0,\epsilon>0.$
Definition 2.2 (Young’s inequality with $\epsilon$ [11]).
Let $1<p,\>q<\infty$, $\frac{1}{p}+\frac{1}{q}=1$. Then
$ab\leq\epsilon a^{p}+C(\epsilon)b^{q},\qquad\qquad a,b>0,\epsilon>0,$
for $C(\epsilon)=(\epsilon p)^{\frac{-q}{p}}q^{-1}$.
Definition 2.3 (Hölder’s inequality [11]).
Assume $1<p,\>q<\infty$, $\frac{1}{p}+\frac{1}{q}=1$. If $u\in{\mathcal{L}^{p}}(\Omega),\>v\in{\mathcal{L}^{q}}(\Omega)$, then we have
$\int_{\Omega}|uv|\mathrm{d}x\leq\|u\|_{{\mathcal{L}^{p}}(\Omega)}\>\|v\|_{{\mathcal{L}^{q}}(\Omega)}.$
Definition 2.4 (Interpolation inequality [11]).
Assume $1\leq s\leq r\leq t\leq\infty$ and $\frac{1}{r}=\frac{\theta}{s}+\frac{(1-\theta)}{t}$, Suppose also $u\in{\mathcal{L}^{s}(\Omega)}\cap{\mathcal{L}^{t}(\Omega)}$. Then $u\in{\mathcal{L}^{r}(\Omega)}$ and
$\big\|u\big\|_{\mathcal{L}^{r}(\Omega)}\leq\big\|u\big\|^{\theta}_{{\mathcal{L}^{s}}(\Omega)}\>\big\|u\big\|^{1-\theta}_{\mathcal{L}^{t}(\Omega)}.$
Definition 2.5 (The Gagliardo-Nirenberg inequality [11]).
[…]
$\big\|D^{j}u\big\|_{{\mathcal{L}^{p}}(\Omega)}\leq C_{1}\big\|D^{m}u\big\|^{\alpha}_{{\mathcal{L}^{r}}(\Omega)}\big\|u\big\|^{1-\alpha}_{{\mathcal{L}^{q}}(\Omega)}+C_{2}\big\|u\big\|_{{\mathcal{L}^{s}}(\Omega)},$
where $s>0$ is arbitrary.
@@
Suppose that $\Omega\subset\mathbb{R}^{n}\>(n\geq 2)$, is a bounded domain with smooth boundary and assume that $q>n$. Moreover, the function $f$ fulfills (1.10) and $k,m\geq 1$, then for each non-negative initial data satisfies (1.9), there exists $T_{\mathrm{max}}\in(0,\infty]$ such that a uniquely determined triple (u,v,w)({u},{v},{w}) of nonnegative functions
@@
The proof can be derived by the standard arguments involving the parabolic regularity theory. For the proof we refer to [19] and [12]. In addition, the non-negativity of the solution in $\Omega\times(0,T_{\mathrm{max}})$ comes from the maximum principle along with (1.9). ∎
@@
(See [3]) For $\int_{\Omega}{u}(x,0)\mathrm{d}x=\int_{\Omega}u_{0}(x)\mathrm{d}x=M_{0}>0$, the mass $\int_{\Omega}{u}(x,t)\mathrm{d}x=M(t)$ satisfies
@@
For each $y_{0}\in{\mathcal{W}^{2,r}}(\Omega)$ such that $\frac{\partial y_{0}(x)}{\partial\nu}=0$ on $\partial\Omega$ and any $g\in{\mathcal{L}^{r}}\big((0,T);{\mathcal{L}^{r}}(\Omega)\big)$, there exists a unique solution
@@
$\int_{0}^{T}\int_{\Omega}|y|^{r}+\int_{0}^{T}\int_{\Omega}|y_{t}|^{r}+\int_{0}^{T}\int_{\Omega}|\Delta y|^{r}\leq C_{r}\int_{0}^{T}\int_{\Omega}|g|^{r}+C_{r}\int_{\Omega}|y_{0}|^{r}+C_{r}\int_{\Omega}|\Delta y_{0}|^{r}.$
@@
$\int_{s_{0}}^{T}\int_{\Omega}e^{sr}|\Delta y|^{r}\leq C^{\prime}_{r}\int_{s_{0}}^{T}\int_{\Omega}e^{sr}|g|^{r}+C^{\prime}_{r}\int_{\Omega}|y(\cdot,s_{0})|^{r}+C^{\prime}_{r}\int_{\Omega}|\Delta y(\cdot,s_{0})|^{r},$
where $C_{r}^{\prime}>0$.
@@
Let $n\geq 2,s_{0}\in(0,T_{\mathrm{max}})$ with $s_{0}<1$ and the initial data ${u}_{0}$, ${v}_{0}$ and ${w}_{0}$ satisfy (1.9) for some $q>n$. Moreover, the function f fulfill (1.10) with $l+k<1+\frac{2}{n}$, where $k\geq 1$ and $1<m<1+\frac{2}{n}$. Suppose that there exists $p>\frac{n}{2}\geq 1$ such that
@@
$\sup_{t>0}\Big(\big\|{u}(\cdot,t)\big\|_{{\mathcal{L}^{\infty}}(\Omega)}+\big\|{v}(\cdot,t)\big\|_{{\mathcal{W}^{1,q}}(\Omega)}+\big\|{w}(\cdot,t)\big\|_{{\mathcal{W}^{1,q}}(\Omega)}\Big)<\infty.$
@@
$\frac{np}{(n-p)_{+}}=\left\{\begin{array}[]{rrll}\hskip-56.9055pt\hfil&&\infty,&\quad\mbox{if}\quad p\geq n,\\ \\ \hskip-56.9055pt\hfil&&\frac{np}{n-p}>n,&\quad\mbox{if}\quad\frac{n}{2}<p<n,\end{array}\right$.
and choose $q<\frac{np}{(n-p)_{+}}$ and $1<q_{0}<q$ fulfilling $n<q_{0}<\frac{np}{(n-p)_{+}}$ which enables to choose $\sigma>1$,   $n<\sigma q_{0}<\frac{np}{(n-p)_{+}}$ and $\sigma q_{0}\leq q$. We fix arbitrary $s_{0}\in(0,T_{\mathrm{max}})$ with $s_{0}<1$ and $t\in(s_{0},T_{\mathrm{max}})$. Applying the variation of constants formula to the second equation of (1.6), we get
${v}(\cdot,t)=e^{-\alpha t}e^{d_{2}t\Delta}{v}_{0}+\int^{t}_{s_{0}}e^{-\alpha(t-s)}e^{d_{2}(t-s)\Delta}f\big(u(\cdot,s)\big)\mathrm{d}s.$
@@
$+K\int^{t}_{s_{0}}e^{-\alpha(t-s)}\Big\|\nabla e^{d_{2}(t-s)\Delta}{u}^{l}(\cdot,s)\Big\|_{{\mathcal{L}^{q_{0}\sigma}(\Omega)}}\mathrm{d}s.$